# Explainable Concept Generation through Vision-Language Preference Learning

### v1

@@ Abstract
Concept-based explanations have become a popular choice for explaining deep neural networks post-hoc because, unlike most other explainable AI techniques, they can be used to test high-level visual “concepts” that are not directly related to feature attributes. For instance, the concept of “stripes” is important to classify an image as a zebra. Concept-based explanation methods, however, require practitioners to guess and collect multiple candidate concept image sets, which can often be imprecise and labor-intensive. Addressing this limitation, in this paper, we frame concept image set creation as an image generation problem. However, since naively using a generative model does not result in meaningful concepts, we devise a reinforcement learning-based preference optimization algorithm that fine-tunes the vision-language generative model from approximate textual descriptions of concepts. Through a series of experiments, we demonstrate the capability of our method to articulate complex, abstract concepts that are otherwise challenging to craft manually. In addition to showing the efficacy and reliability of our method, we show how our method can be used as a diagnostic tool for analyzing neural networks.
@@ 1 Introduction
Humans utilize high-level concepts as a medium for providing and perceiving explanations. In this light, post-hoc concept-based explanation techniques, such as Testing with Concept Activation Vectors (TCAV) [1], have gained great popularity in recent years. Their ability to use abstractions that are not necessarily feature attributes or some pixels in test images helps with communicating these high-level concepts with humans. For instance, as demonstrated in TCAV, the concept of stripes is important to explain why an image is classified as a zebra, whereas the concept of spots is important to explain why an image is classified as a jaguar. Given 1) a set of such high-level of concepts, represented as sample images (e.g., a collection of stripe images and a collection of spot images) and 2) test images of the class (i,e., zebra images), TCAV assigns a score to each concept on how well the concept explains the class decision (i.e., zebra).
Although concept-based XAI methods are a good representation, their requirement to create collections of candidate concept sets necessitate the human to know which concepts to test for. This
Figure 1: Our algorithm, RLPO, gradually updates stable diffusion weights (black arrows) while selecting the potential text prompts (a). Each step in this update process provides an explanation at a different level of abstraction.
@@
is typically done by guessing what concepts might matter and manually extracting such candidate concept tests from existing datasets. While the stripe-zebra analogy is attractive as an example, where it is obvious that stripes is important to predict zebras, in most applications, we cannot guess what concepts to test for, limiting the usefulness of concept-based methods in testing real-world systems. Additionally, even if a human can guess a few concepts, it does not encompass most concepts a DNN has learned because the DNN was trained without any human intervention. Therefore, it is important to automatically find human-centric concepts that matter to the DNN’s decision-making process. As attempts to automatically discover and create such concept sets, several work has focused on segmenting the image and using these segments as potential concepts, either directly [2] or through factor analysis [3, 4]. In such methods, which we call as retrieval methods, because the extracted concept set is already part of the test images, it is difficult for them to imagine new concepts that do not have a direct pixel-level resemblance to the original image class. For instance, it is arguable if some patches of zebra—instead of stripes—qualify as high-level concepts to explain the zebra class.
@@
By departing from existing concept set creation practices of human handcrafting and retrieval, we redefine concept set creation as a concept generation problem. Modern generative models such as stable diffusion (SD) can produce noise-free, realistic images. Nevertheless, since a generative model can generate arbitrary images, we need to guide it to produce what we desire by using text prompts. One obvious approach is to engineer long, descriptive text prompts to generate concepts. However, engineering such prompts is not realistic. Therefore, to automate prompting, we extract keywords related to the image using an image-to-text model. We propose a reinforcement learning-based preference optimization (RLPO) algorithm that guides the generative model to automatically generate meaningful concepts based on these seed keyword prompts. Preferences are solely decided by the explanation score—not by a human—that the deep RL (DRL) algorithm is trying to optimize.
By formulating concept discovery as a concept generation cum search problem offers numerous advantages. First, we can find concepts that truly matters to the neural network. Some of these concepts would not pop up in humans head until they see them. Also, unlike existing retrieval methods, the proposed method can generate concepts that are not part of test images (Fig. 3). Further, because of how we use preference optimization to gradually make the SD process closer to the class image from a high-level concept, for the first time in concept-based XAI, we can generate concepts with different abstraction levels. Architecturally, the modularity of our approach allows for the easy replacement of the generative model and other components with any future state-of-the-art models.
@@ 2 Preliminaries and Related Work
Testing with Concept Activation Vectors (TCAV): The TCAV score quantifies the importance of a “concept” for a specific class in a DNN classifier [1]. Here, a concept is defined broadly as a high-level, human-interpretable idea such as stripes, sad faces, etc. A concept (e.g., stripes), $c$, is represented by sample images, $X_{c}$ (e.g., images of stripes). In TCAV, a human has manually collected these sample concept images based on educated guesses, whereas our objective is to automatically generate them. For a given set of test images, $X_{m}$ (e.g., zebra images), that belongs to the same decision class (e.g., zebra), $m$, TCAV is defined as the fraction of test images for which the model’s prediction increases in the direction of the concept. By decomposing the DNN under test as $f(x)=f_{2}(f_{1}(x))$, where $f_{1}(x)$ is the activation at layer $l$, TCAV score is computed as,
@@
$S_{c,m}=\frac{1}{|X_{m}|}\sum_{X_{m}}\mathbb{I}\left(\frac{\partial\text{output}}{\partial\text{activations}}\cdot{(c\enspace\text{direction})}>0\right)=\frac{1}{|X_{m}|}\sum_{x_{i}\in X_{m}}\mathbb{I}\left(\frac{\partial f(x_{i})}{\partial f_{1}(x_{i})}\cdot v>0\right)$ (1)
@@
Here, $\mathbb{I}$ is the indicator function that counts how often the directional derivative is positive. Concept activations vector (CAV), $v$, is the normal vector to the hyperplane that separates activations of concept images, $\{f_{1}(x);x\in X_{c}\}$, from activations of random images, $\{f_{1}(x);x\in X_{r}\}$. Refer to supplementary section 2.1 for details on the TCAV settings.
@@
ACE [2] introduced a way to automatically find concepts by extracting relevant concepts from the input class. It uses image segmentation of multiple sizes to get a pool of segments and then grouped them based on similarity to compute TCAV scores. Though the ACE concepts are human understandable, they are very noisy. EAC [5] extracts concepts through segmentation. CRAFT [3] introduced a recursive strategy to detect and decompose concepts across layers. Lens [4] elegantly unified concept extraction and importance estimation as a dictionary learning problem. However, since all these methods obtain concepts from test images, the concepts they generate tend to be very similar to the actual class, making it challenging to maintain the “high-levelness” of concepts. In contrast, we generate concepts from a generative model. Under generative models, LCDA [6] simply queries an LLM to get attributes but does not generate concepts.
@@
Deep Q Networks (DQN): DQN [7] is a DRL algorithm that combines Q-learning with deep neural networks. It is designed to learn optimal policies in environments with large state and action spaces by approximating the Q-value function using a neural network. A separate target network, $Q_{\text{target}}(s,a^{\prime},\theta^{\prime})$, Here $a^{\prime}$ is $argmaxQ(s_{next},a)$ which is a copy of the Q-network with parameters $\theta^{\prime}$, is updated less frequently to provide stable targets for Q-value updates,
@@
Here, $s_{t}$ is the state at step $t$, $a_{t}$ is the action taken in state $s_{t}$, and $r_{t}$ is the reward received after taking action $a_{t}$. The parameters $\alpha$ and $\gamma$ are learning rate and discount factor, respectively. DQNs are used for controlling robots [8, 9, 10], detecting failures [11], etc.
Direct Preference Optimization (DPO): DPO [12] is a technique used to ensure models, such as large language models, learn to align its outputs with human preference by asking a human which of its generated output is preferred. In this study, preference is provided by the TCAV score instead of a human, and it is used to align a text-to-image generative model that generates concepts with the neural network under test.
3 Methodology: Explainable Concept Generation
1:Run pre-processing and get the action space
2:for each episode do
3:     for each time step $t$ do
4:         Execute $a_{t}$ and pick a keyword
5:         Generate image groups $G_{1}$ $G_{2}$
6:         Evaluate TCAV scores $S_{1}$ $S_{2}$
7:         Update SD based on better score
8:         Compute reward
9:     end for
@@ 3 Methodology: Explainable Concept Generation
Algorithm 1 The RLPO algorithm.
Our objective is generating concept images, $\mathcal{C}$, that provide a higher TCAV score, $S_{c,m}$. To this end, we leverage the state-of-the-art text-to-image generative models to generate high quality explainable concepts. However, because the search space of potential text prompts is too large, we use DRL to search for text that are relevant. Our algorithm, named reinforcement learning-based preference optimization (RLPO), picks a potential keyword using RL and optimizes stable diffusion weights to generate images that have a preference for higher TCAV scores. As summarized in Fig. 1 and Algorithm 1, we describe this process in the rest of this section. This work help engineers and other users of DNNs to obtain high-level visual explanations of under which conditions the neurons are activated in a DNN, making it a useful diagnostic tool.
Notation: Our framework contains multiple machine learning models. First, we have a pre-trained neural network classifier that we want to explain. We call this the network under test (NUT), which is indicated by $f(\cdot)$. We then have a generative neural network, $g(\cdot)$, whose purpose is generating concept image sets, given some text prompts. In this paper, we use Stable Diffusion (SD) v1-5 as the generator as it is a state-of-the-art generative model that can generate realistic images. If the weights of the SD model are $w$, for a small constant $\lambda$, we augment it as $w+\lambda ab$, where $A$ and $B$ are low-rank matrices that we fine-tune using DPO [12]. The core search algorithm that we train contains a DQN which is denoted by $h(\cdot)$. We use off-the-shelf visual question answering (VQA) to generate the action space for the DQN. There is also a linear classification model that we train to extract explainable concept directions using TCAV.
@@ 3.1 The Rationales Behind Design Choices
Before presenting the algorithm in detail, we provide the rationale for design choices, which are validated through ablation studies and comparisons in the Experiments section.
Rationale 1: Why concept generation is a better idea. Let us denote the set of human-interpretable concepts that the NUT has learned be $\mathcal{C}_{N}$. If we use concept-based explanation the traditional way [1, 13], then the end users need to manually guess what concepts to test for. Automatically retrieving the concept set by segmenting test images [5] also results in a limited concept set. In contrast, a SOTA generative model can generate high quality images.
@@
As the saying goes, “a picture is worth a thousand words,” it is much easier for people to explain and understand high-level concepts when images are used instead of language. For instance, we need a long textual description such as “The circles are centered around a common point, with alternating red and white colors creating a pattern” to describe a simple image of a dart board (i.e., Target Co. logo). Therefore, we keep our ultimate concept representation as images. However, controlling a generative model from visual inputs is much harder. However, since human language can be used as a directed and easier way to seed our thought process, as the saying goes, “words flow easier than paint,” we control the use of text prompts. Since the vastness of the search space cannot be handled by most traditional search strategies, we resort to a DQN for controlling text. Since simple text alone cannot generate complex, high-level visual concepts, in each DQN update step, we use preference optimization to further guide the search process to guide towards more preferred outcome, allowing the DQN to focus on states similar to the target. This approach improves our starting points for each DQN episode, enabling more efficient search and incremental progress towards the desired target.
3.2 The Deep Reinforcement Learning Formulation
Our objective of using DRL is automatically controlling text prompts used in the text-to-image generative model. As text prompts, we start with a list of keywords, $\mathcal{K}$, that have the potential to generate meaningful concept images after many DRL episodes. How we extract these keywords automatically will be discussed in detail in Section 3.3. We would like to highlight that these keywords act as seeds and they do not directly impact the accuracy of the DRL algorithm although good seeds will make the policy learning faster. We setup our RL state-action at iteration $t$ as follows:
  • Action $a_{t}$: Selecting a keyword, $k_{t}\in\mathcal{K}$, that best influences concept image generation
  • State $s_{t}$: Concept images generated from the keyword prompt, $k_{t-1}$.
Our objective in DRL is to learn a policy, $\pi:s\to a$, that takes actions (i.e., picking keywords) leading to explainable states (i.e., correct concept images) from proxy states (i.e., somewhat correct concept images). This traversal is illustrated in Fig. 1 and formally defined below.

### v2

@@ Abstract
Concept-based explanations have become a popular choice for explaining deep neural networks post-hoc because, unlike most other explainable AI techniques, they can be used to test high-level visual “concepts” that are not directly related to feature attributes. For instance, the concept of “stripes” is important to classify an image as a zebra. Concept-based explanation methods, however, require practitioners to guess and collect multiple candidate concept image sets, which can often be imprecise and labor-intensive. Addressing this limitation, in this paper, we frame concept image set creation as an image generation problem. However, since naively using a generative model does not result in meaningful concepts, we devise a reinforcement learning-based preference optimization (RLPO) algorithm that fine-tunes the vision-language generative model from approximate textual descriptions of concepts. Through a series of experiments, we demonstrate the capability of our method to articulate complex and abstract concepts which aligns with the test class that are otherwise challenging to craft manually. In addition to showing the efficacy and reliability of our method, we show how our method can be used as a diagnostic tool for analyzing neural networks.
@@ 1 Introduction
Humans utilize high-level concepts as a medium for providing and perceiving explanations. In this light, post-hoc concept-based explanation techniques, such as Testing with Concept Activation Vectors (TCAV) (Kim et al., 2018), have gained great popularity in recent years. Their ability to use abstractions that are not necessarily feature attributes or some pixels in test images helps with communicating these high-level concepts with humans. For instance, as demonstrated in TCAV, the concept of stripes is important to explain why an image is classified as a zebra, whereas the concept of spots is important to explain why an image is classified as a jaguar. Given 1) a set of such high-level concepts, represented as sample images (e.g., a collection of stripe images and a collection of spot images) and 2) test images of the class (e.g., zebra images), TCAV assigns a score to each concept on how well the concept explains the class decision (i.e., zebra).
Figure 1: (a) Our proposed algorithm, RLPO, iteratively refines the concepts $c_{i}$ that can be generated by a Stable Diffusion (SD) model by optimizing SD weights based on an action $a_{i}$. Each step in this update process provides an explanation at a different level of abstraction. (b) Three concepts identified by our approach for the zebra class. Concepts are represented as images generated by SD.
@@
Although concept-based XAI methods are a good representation, their requirement to create collections of candidate concept sets necessitate the human to know which concepts to test for. This is typically done by guessing what concepts might matter and manually extracting such candidate concept tests from existing datasets. While the stripe-zebra analogy is attractive as an example, where it is obvious that stripes is important to predict zebras, in most applications, we cannot guess what concepts to test for, limiting the usefulness of concept-based methods in testing real-world systems. Additionally, even if a human can guess a few concepts, it does not encompass most concepts a DNN has learned because the DNN was trained without any human intervention. Therefore, it is important to automatically find human-centric concepts that matter to the DNN’s decision-making process.
As attempts to automatically discover and create such concept sets, several work has focused on segmenting the image and using these segments as potential concepts, either directly (Ghorbani et al., 2019) or through factor analysis (Fel et al., 2023; 2024). In such methods, which we refer to as retrieval methods, because the extracted concept set is already part of the test images as shown in Fig. 2 (Retrieved concepts), it is difficult for them to imagine concepts that do not have a direct pixel-level resemblance to the original image class. For instance, it is more likely that such methods provide patches of zebra as concepts instead of stripes.
@@
By departing from existing concept set creation practices of human handcrafting and retrieval, we redefine concept set creation as a concept generation problem. Modern generative models such as stable diffusion (SD) can be used for generating realistic images. Nevertheless, since a generative mode generates arbitrary images, we need to carefully guide it to generate explanatory images. One obvious approach is to engineer long, descriptive text prompts to generate concepts. However, engineering such prompts is not realistic. Therefore, to automate this process, as shown in Fig. 1, we propose a method, named reinforcement learning-based preference optimization (RLPO), to update SD weights. At its core, we devised a deep reinforcement learning algorithm gradually update SD weights to generate concept images that have a higher explanation score. The contributions of this paper can be summarized as follows:
  1. 1.
    We propose a method, named RLPO, to “generate” concepts that truly matters to the neural network. Some of these concepts would not pop up in humans head until they see them. Also, unlike existing retrieval methods, RLPO can generate concepts that are not part of test images (Fig. 2).
  2. 2.
    Because of how we use model’s preference to gradually make the SD process closer to the target class from a high-level concept, for the first time in concept-based XAI, we can generate concepts with different abstraction levels.
  3. 3.
    In addition to demonstrating the novelty, abstractness, and diversity of generated concepts, we experimentally verify their generalizability using an NLP sentiment analysis tasks and the actionability by leveraging concepts as a tool for fine-tuning.
@@ 2 Preliminaries and Related Work
Testing with Concept Activation Vectors (TCAV): The TCAV score quantifies the importance of a “concept” for a specific class in a DNN classifier (Kim et al., 2018). Here, a concept is defined broadly as a high-level, human-interpretable idea such as stripes, sad faces, etc. A concept (e.g., stripes), $c$, is represented by sample images, $X_{c}$ (e.g., images of stripes). In TCAV, a human has manually collected these sample concept images based on educated guesses, whereas our objective is to automatically generate them. For a given set of test images, $X_{m}$ (e.g., zebra images), that belongs to the same decision class (e.g., zebra), $m$, TCAV is defined as the fraction of test images for which the model’s prediction increases in the “direction of the concept.” By decomposing the DNN under test as $f(x)=f_{2}(f_{1}(x))$, where $f_{1}(x)$ is the activation at layer $l$, TCAV score is computed as,
@@
$TS_{c,m}=\frac{1}{|X_{m}|}\sum_{X_{m}}\mathbb{I}\left(\frac{\partial\text{output}}{\partial\text{activations}}\cdot{(c\enspace\text{direction})}>0\right)=\frac{1}{|X_{m}|}\sum_{x_{i}\in X_{m}}\mathbb{I}\left(\frac{\partial f(x_{i})}{\partial f_{1}(x_{i})}\cdot v>0\right)$ (1)
@@
Here, $\mathbb{I}$ is the indicator function that counts how often the directional derivative is positive. Concept activations vector (CAV), $v$, is the normal vector to the hyperplane that separates activations of concept images, $\{f_{1}(x);x\in X_{c}\}$, from activations of random images, $\{f_{1}(x);x\in X_{r}\}$. Refer to Appendix C.1 for details on the TCAV settings.
@@
ACE (Ghorbani et al., 2019) introduced a way to automatically find concepts by extracting relevant concepts from the input class. It used segmentation over different resolution to get a pool of segments and then grouped them based on similarity to compute TCAV scores. Though the ACE concepts are human understandable, they are noisy because of the segmentation and clustering errors. As a different method, EAC (Sun et al., 2024) extracts concepts through segmentation. CRAFT (Fel et al., 2023) introduced a recursive strategy to detect and decompose concepts across layers. Lens (Fel et al., 2024) elegantly unified concept extraction and importance estimation as a dictionary learning problem. However, since all these methods obtain concepts from test images, the concepts they generate tend to be very similar to the actual class (e.g., a patch of zebra as a concept to explain the zebra instead of stripes as a concept), making it challenging to maintain the “high-level abstractness” of concepts. In contrast, we generate concepts from a generative model. Under generative models, LCDA (Yan et al., 2023) simply queries an LLM to get attributes but does not generate concepts.
@@
Deep Q Networks (DQN): DQN (Mnih et al., 2015) is a deep RL algorithm that combines Q-learning with deep neural networks. It is designed to learn optimal policies in environments with large state and action spaces by approximating the Q-value function using a neural network. A separate target network, $Q_{\text{target}}(s,a^{\prime},\theta^{\prime})$, Here $a^{\prime}$ is $\arg\max Q(s_{next},a)$ which is a copy of the Q-network with parameters $\theta^{\prime}$, is updated less frequently to provide stable targets for Q-value updates,
@@
Here, $s_{t}$ is the state at step $t$, $a_{t}$ is the action taken in state $s_{t}$, and $r_{t}$ is the reward received after taking action $a_{t}$. The parameters $\alpha$ and $\gamma$ are learning rate and discount factor, respectively.
Preference Optimization: Directly optimizing generative models with preference data was first introduced in Direct Preference Optimization (DPO) (Rafailov et al., 2024). It is a technique used to ensure models, such as large language models, learn to align its outputs with human preference by asking a human which of its generated output is preferred. This technique was later extended to diffusion models in Diffusion-DPO (Wallace et al., 2023), where they updated Stable Diffusion XL model using Pick-a-Pic dataset (human preferred generated image dataset). Unlike traditional image or text generation tasks, where the dataset for human preferred outputs are readily available, it is hard to have a general enough dataset for XAI tasks. To counter this problem, we provide preference information by using the TCAV score instead of a human, and use it to align the text-to-image generative model to generate concept images that matters for the neural network under test.
3 Methodology: Reinforcement Learning-based Preference Optimization
Figure 3: Overview of the RLPO framework with its dynamic environment interaction. The RL policy selects actions (seed prompts) which generates concept sets (G1, G2) scored through TCAV. Reward is calculated based on the scores obtained for both the sets. Simultaneously, best set is determined based on the scores obtained, which is used to update the LoRA layer of the SD model.
Our objective is to find a set of concept images, $\mathcal{C}$, that maximize the TCAV scores, $TS_{c,m}$. To this end, we leverage the state-of-the-art text-to-image generative models to generate high quality explainable concepts. However, because the search space of potential text prompts is too large, we use deep RL to guide the image generation process. As described in Fig. 3 and Algorithm 1, our algorithm, RLPO, picks potential keywords from an automatically generated list of keywords using RL and optimizes stable diffusion weights to generate images that have a preference for higher TCAV scores. This process is described in the rest of this section.
Algorithm 1 The RLPO algorithm. Appendix C.2 for the expanded algorithm.
1:Input: Set of test images, $f(\cdot)$
2:Run pre-processing and get the seed prompts (action space)
3:for each episode do
4:     for each time step $t$ do
5:         Execute $a_{t}$ and pick a keyword
6:         Generate image groups $G_{1}$ $G_{2}$
7:         Evaluate TCAV scores $TS_{1}$ $TS_{2}$
8:         Update SD based on better score
9:         Compute reward
@@ 3 Methodology: Reinforcement Learning-based Preference Optimization
11:end for
12:Output: Set of concept images
Notation: Our framework contains three core deep learning models: the network under test $f(\cdot)$, the image generator $g(\cdot)$, and the deep RL network $h(\cdot)$. First, we have a pre-trained neural network classifier that we want to explain. We then have a generative neural network, whose purpose is generating concept image sets, given some text prompts. In this paper, we use Stable Diffusion (SD) v1-5 as the generator as it is a state-of-the-art generative model that can generate realistic images. If the weights of the SD model are $w$, for a small constant $\lambda$, we augment it as $w+\lambda ab$, where $A$ and $B$ are low-rank matrices that we fine-tune using preference optimization (Rafailov et al., 2024). The core search algorithm that we train is a DQN.
@@ 3.1 The Rationales Behind Design Choices
Before presenting the algorithm in detail, we provide the rationale for design choices, which are validated through ablation studies in Section 4.1-4.2, and comparisons in the Section 4.3-4.4.
Rationale 1: Why concept generation is a better idea. Let us denote the set of human-interpretable concepts that the $f(\cdot)$ has learned be $\mathcal{C}_{N}$. If we use concept-based explanation the traditional way (Kim et al., 2018; Schut et al., 2023), then the end users need to manually guess what concepts to test for. Automatically retrieving the concept set by segmenting test images (Sun et al., 2024) also results in a limited concept set. In contrast, a SOTA generative model can generate high quality images. We provide more theoretical insights in Appendix B.
@@
As the saying goes, “a picture is worth a thousand words,” it is much easier for people to explain and understand high-level concepts when images are used instead of language. For instance, we need a long textual description such as “The circles are centered around a common point, with alternating red and white colors creating a pattern” to describe a simple image of a dart board (i.e., Target Co. logo). Therefore, we keep our ultimate concept representation as images. However, controlling a generative model from visual inputs is much harder. However, since human language can be used as a directed and easier way to seed our thought process, as the saying goes, “words flow easier than paint,” we control the use of text prompts. Since the vastness of the search space cannot be handled by most traditional search strategies, we resort to a DQN for controlling text. Since simple text alone cannot generate complex, high-level visual concepts, in each DQN update step, we use preference optimization to further guide the search process towards more preferred outcome, allowing the DQN to focus on states similar to the target. This approach improves our starting points for each DQN episode, enabling more efficient search and incremental progress towards the desired target.
3.2 Extracting Seed Prompts
Since a generative model can generate arbitrary images, if we provide good starting point for optimization then the convergence to explainable states would be faster. In this paper, to extract seed prompts for a particular class we make use of the off-the-shelf VQA model followed by several preprocessing steps, as described in Appendix C.3. We also explore how random gibberish prompts can be used as seed prompts in Appendix C.4 which did not yield useful concepts.
3.3 Deep Reinforcement Learning Formulation
Our objective of using deep RL is automatically controlling text input of Stable Diffusion. As text input, we start with seed prompts from Section 3.2, $\mathcal{K}$, that have the potential to generate meaningful concept images after many deep RL episodes. We setup our RL state-action at iteration $t$ as,
  • Action $a_{t}$: Selecting a seed prompt, $k_{t}\in\mathcal{K}$, that best influences concept image generation.
  • State $s_{t}$: Preferred concept images generated from the seed prompt, $k_{t-1}$.
  • Reward $r_{t}$: Reward $r_{t}$ is proportional to the TCAV score computed at state $s_{t}$ on action $a_{t}$, adjusted by a monotonically increasing scaling factor $\xi_{t,k}$. As each seed concept reaches the explainable state at different times, this factor is introduced to scale the reward over time $t$ for each unique seed concept $k$. Since the $g(.)$ is getting optimized at each time step $t$. The scaling factor is updated as $\xi_{t+1,k}\leftarrow\min(1,\frac{\xi_{t,k}+1}{T})$, where T is total number of RL steps. Therefore, the expected cumulative adjusted reward is $R(\pi)=\mathbb{E}[\sum_{t=0}^{T}\xi_{t}\cdot r_{t}(s_{t},a_{t})]$.
Our objective in deep RL is to learn a policy, $\pi:s\to a$, that takes actions (i.e., picking a seed prompt) leading to explainable states (i.e., correct concept images) from proxy states (i.e., somewhat correct concept images). We formally define explainable state and proxy state as follow: